# Anomalous Energy Injection in Turbulent Neutron Star Cores

### v1

@@ Abstract
Neutron star glitches—sudden increases in rotational frequency—are thought to result from angular momentum transfer via quantized vortices in the superfluid core. Using a two-dimensional rotating atomic Bose-Einstein condensate with phenomenological damping and a pinning potential to mimic the crust, we model the dynamics underlying these events. Our simulations reveal a transient Kolmogorov-like turbulent cascade ($k^{-5/3}$) that transitions to a Vinen-like scaling ($k^{-1}$). We identify an anomalous secondary injection mechanism driven primarily by quantum pressure, which sustains turbulent fluctuations in pulsar glitches. By tuning the damping coefficient $\gamma$, we determine an optimal regime for energy transfer. These findings provide a robust analogy for neutron star glitch phenomena and offer new insights into turbulence in extreme astrophysical environments.
@@
Introduction: Decaying neutron stars exhibit sudden increases in rotation frequency, known as glitches [1], caused by the transfer of angular momentum from quantized vortices in the neutron superfluid to the outer crust of the pulsar [2, 3]. This transfer is primarily triggered by the vortex avalanche mechanism, in which quantum vortices become pinned to crustal nuclei [4], preventing the superfluid from spinning down at the same rate as the crust. The resulting rotational lag between the crust and superfluid generates a Magnus force on the vortices [5]. Once a critical velocity is reached, the quantized vortices unpin from their pinning sites [6]. These vortices are then expelled toward the crust, transferring their angular momentum and producing a glitch [7]. This glitch process involves a complex vortex avalanche encompassing millions of vortices, triggering their depinning [8].
@@
Modeling the depinning avalanche mechanism poses significant challenges. Vortex depinning initiates when a vortex, initially trapped by a pinning site, overcomes the attractive force and begins moving in the direction of the applied superfluid flow. Several factors, such as the critical velocity of the superfluid flow and the size and shape of the pinning site, affect the overall dynamics of depinning [9]. However, an alternative analysis suggests that glitches emerge due to a quantum transition in which the core transforms into a superconducting gluon-quark superfluid state, with resultant vortex depinning occurring at the boundary layer between the core and the surrounding dissipative medium [3]. In this letter, we demonstrate the presence of anomalous energy transfer from quantum pressure to incompressible kinetic energy in the turbulent state of the neutron star core, which sustains pulsar glitches.
@@
In recent years, rotating Bose-Einstein condensates (BECs) have been widely employed to model the dynamics of quantized vortices in different astrophysical systems, such as neutron superfluids in strong gravitational wells [10, 11, 12], turbulent dark matter [13], pulsar glitches [14]. Verma et al. [15] and Shukla et al. [14] developed a minimal model for the emergence of the superfluid glitches based on the interaction of the neutron-superfluid vortices and proton-superconductor flux tubes. The inherent compressibility of the BECs renders them susceptible to quantum turbulence (QT) under dynamic instabilities. This turbulence is primarily characterized by the breakdown of vortices in a self-similar process, manifesting as a power-law scaling of $k^{-5/3}$ [16, 17], known as the Kolmogorov spectrum. In addition to the Kolmogorov regime, QT in BECs also exhibits Vinen turbulence, characterized by random vortex distributions in superfluids, indicated by a $k^{-1}$ scaling. Novel turbulence regimes have been identified in recent years, including strong quantum turbulence [18, 19], marked by density fluctuations and non-polarized vortex lines; rotational quantum turbulence [20, 21, 22, 23]; and turbulence in self-gravitating dark-matter BEC candidates [24, 13].
@@
Previous studies have explored the nature of glitches and vortex avalanches in the context of pinning potentials and superfluid behavior. However, the role of superfluid-crust interactions in inducing turbulent flow remains largely unexamined. In this letter, we investigate the characteristics of turbulence and associated velocity flows arising from such interactions. Furthermore, it analyzes the mechanisms driving this turbulence, facilitated by pinning sites and the dynamic spin-down of the condensate. The damping effects on the system are also evaluated, identifying an optimal damping coefficient that maximizes turbulence strength. Recent studies indicate that quantum turbulence suppresses collective excitation modes in BECs [25, 26]. As these modes drive vortex avalanches and glitch phenomena, quantifying the onset of turbulence in such systems is essential.
@@
where $\psi\equiv\psi(\mathbf{r},t)$ is the condensate wave function, with $\mathbf{r}\equiv(x,y)$. $\nabla^{2}=\partial_{x}^{2}\,+\partial_{y}^{2}$ represents the two-dimensional Laplacian, $\gamma$ corresponds to the parameterized damping coefficient, and $g$ is the nonlinear interaction strength given by $g=800$. The time-dependent rotation frequency $\Omega(t)$ describes the rotational deceleration profile of the condensate as follows:
@@
where $\Omega_{0}$ is the initial rotation frequency of the condensate, and $t_{s}$ is the spin-down time. The $z$-component of the angular momentum is given by $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$.
@@
where $\tilde{V}(x_{\theta},y_{\theta})=3+\cos(n_{\mathrm{crust}}x_{\theta})+\cos(n_{\mathrm{crust}}y_{\theta})$, and the rotated coordinates are $x_{\theta}=\cos(\theta(t))x_{p}+\sin(\theta(t))y_{p}$ and $y_{\theta}=-\sin(\theta(t))x_{p}+\cos(\theta(t))y_{p}$. Here, $n_{\mathrm{crust}}$ determines the number of pinning sites, $r_{\mathrm{crust}}$ is the radius at which $V_{\mathrm{crust}}$ is maximized, and $\Delta r_{\mathrm{crust}}$ represents the crust thickness. The crust potential parameters are based on Refs. [14, 15]. The centrifugal potential, $V_{\mathrm{cent}}(\mathbf{r},t)=\frac{1}{2}\Omega(t)^{2}r^{2}$, ensures uniform condensate density under rotation. Following Refs. [27, 28, 29], we adopt a coherent length scale of $\xi=10~{}\mathrm{fm}$, yielding an energy scale of $\epsilon=207~{}\mathrm{keV}$. With these parameters, the time scale is calculated as $\tau=3.2\times 10^{-21}~{}\mathrm{s}$. We compute the kinetic energy components and their spectra based on the analysis of their numerical implementation in Ref. [30]. The simulation involves numerically solving the Gross-Pitaevskii equation using the split-step Crank-Nicolson method [31, 32] in a computational domain of size $512\times 512$, with spatial steps $dx=dy=0.05$ and a time step $dt=10^{-3}$ to ensure numerical stability and convergence. The numerical simulations are performed on an NVIDIA A100 GPU using CUDA C codes developed based on [33]. Initially, we prepare the condensate and crust with an initial rotation frequency $\Omega_{0}=2$ through imaginary-time iterations ($t\to-\mathrm{i}t$). This converged profile is then spun down at varying rates, characterized by the spin-down time $t_{s}$, in real-time. We first consider a case with a static crust and no dissipation ($\gamma=0$) while exploring the effects of damping in subsequent analyses.
@@
Further, the vortex avalanche triggered by Magnus flow results in an instantaneous recoupling of the superfluid and normal components, accompanied by a rapid transfer of angular momentum. This abrupt transition in superfluid systems may induce turbulence but cannot be confirmed using density profiles alone. To characterize the turbulent flow in the system, we therefore analyze the incompressible and compressible kinetic energy spectra of the condensate.
@@
As shown in Fig. 3, the vortex density exhibits a distinct scaling behavior with respect to time, further confirming the turbulent behavior of the condensate. During the initial Kolmogorov turbulent regime, the vortex-line density $l_{v}$ scales as $t^{-3/2}$, transitioning to $t^{-1}$ scaling in the later Vinen turbulent regime. This temporal decay of vortex-line density is attributed to vortex breakdown in the Kolmogorov regime, as well as to sound radiation (Kelvin-wave emission) and vortex reconnection in the Vinen regime [34, 35, 36]. For shorter spin-down times, the $t^{-3/2}$ scaling persists over a longer timescale, corresponding to a broader spatial extent of the $k^{-3/2}$ scaling in the incompressible kinetic energy spectrum (Fig. 2).
@@
Although glitch events are explained by the transfer of angular momentum [37], analyzing the onset of turbulence and the dynamics of quantized vortices is required to fully understand the transport of kinetic energy components, which, in turn, enables a new perspective on vortex depinning. The evolution of net energy transport between the kinetic components, shown in Fig. 4, provides critical insight into the post-spin-down dynamics and explains the persistence of vortex flow. The energy exchange between the compressible–quantum pressure and incompressible–compressible components does not significantly influence the turbulent dynamics. However, the energy exchange between the incompressible and quantum pressure components, denoted $E_{\mathrm{ex}}^{iq}$, is initially negative, indicating energy transfer from the incompressible to the quantum pressure component. After the condensate spins down, $E_{\mathrm{ex}}^{iq}\to 0$, as no further incompressible energy is transferred to quantum pressure. For sufficiently short spin-down times [see Fig. 4(a)–(b)], the exchange energy becomes positive, indicating an injection of incompressible energy from the quantum pressure component. This secondary injection mechanism sustains turbulent behavior in the condensate, even after the loss of rotational forcing, which explains the existence of vortex flow after the spin-down time is reached. For longer spin-down times, the peak of the secondary injection diminishes, reducing the intensity of turbulence.
@@
Turbulence significantly influences pulsar glitch phenomena via spin-down profiles and dissipative interactions, providing essential insights into the internal dynamics of neutron stars. This study is particularly valuable for exploring inhomogeneous two-dimensional quantum turbulence and its associated dynamics, especially the emergence of a secondary energy injection mechanism.
@@ Acknowledgements.
The work of P.M. is supported by the Ministry of Education-Rashtriya Uchchatar Shiksha Abhiyan (MoE RUSA 2.0): Bharathidasan University – Physical Sciences.

### v2

# Anomalous Energy Injection in the Gross-Pitaevskii Framework for Turbulence in Neutron Star Glitches

@@ Abstract
Neutron star glitches—sudden increases in rotational frequency—are thought to result from angular momentum transfer via quantized vortices in the superfluid core. To investigate the underlying superfluid dynamics, we employ a two-dimensional rotating atomic Bose-Einstein condensate described by a damped Gross-Pitaevskii equation with an imposed pinning potential that serves as a simplified analogue of a crust. Within this minimal framework, we examine the emergence and evolution of turbulent vortex motion following impulsive perturbations reminiscent of glitch-like forcing. Our simulations reveal a transient Kolmogorov-like turbulent cascade ($k^{-5/3}$) that transitions to a Vinen-like scaling ($k^{-1}$). We identify an anomalous secondary injection mechanism driven primarily by quantum pressure, which can sustain turbulent fluctuations in such a system. By tuning the damping coefficient $\gamma$, we determine an optimal regime for energy transfer. While idealized, these findings illustrate how quantum turbulence with multiple scaling regimes can arise in pinned, rotating superfluids, and they suggest possible qualitative connections to vortex-mediated dynamics in neutron stars and other astrophysical superfluid systems.
@@
Introduction: Decaying neutron stars exhibit sudden increases in rotation frequency, known as glitches [1], caused by the transfer of angular momentum from quantized vortices in the neutron superfluid to the outer crust of the pulsar [2, 3]. This transfer is primarily triggered by the vortex avalanche mechanism, in which quantum vortices become pinned to crustal nuclei [4], preventing the superfluid from spinning down at the same rate as the crust. The resulting rotational lag between the crust and superfluid generates a Magnus force on the vortices [5]. Once a critical velocity is reached, the quantized vortices unpin from their pinning sites [6, 7]. These vortices are then expelled toward the crust, transferring their angular momentum and producing a glitch [8]. This glitch process involves a complex vortex avalanche encompassing millions of vortices, triggering their depinning [9].
@@
Modeling the depinning avalanche mechanism poses significant challenges. Vortex depinning initiates when a vortex, initially trapped by a pinning site, overcomes the attractive force and begins moving in the direction of the applied superfluid flow. Several factors, such as the critical velocity of the superfluid flow and the size and shape of the pinning site, affect the overall dynamics of depinning [10]. However, an alternative analysis suggests that glitches emerge due to a quantum transition in which the core transforms into a superconducting gluon-quark superfluid state, with resultant vortex depinning occurring at the boundary layer between the core and the surrounding dissipative medium [3]. In this Letter, we demonstrate anomalous energy transfer from quantum pressure to incompressible kinetic energy, thereby proposing a candidate mechanism for turbulent dynamics within a Gross-Pitaevskii analog of neutron star glitches.
Rotating Bose-Einstein condensates (BECs) provide a quantitatively precise, experimentally tunable, and microscopically resolvable analogue for investigating vortex avalanche processes in neutron-star superfluids [11, 12, 13]. Verma et al. [14] and Shukla et al. [15] developed a minimal model for the emergence of the superfluid glitches, based on the interaction of the neutron-superfluid vortices and proton-superconductor flux tubes. More recently, Poli et al. [16] extended Gross–Pitaevskii models of pinned superfluids by including dipolar interactions and realistic crustal pinning potentials, demonstrating the emergence of self-organised criticality in simulated glitch events. Although these and related Gross–Pitaevskii–based approaches have successfully reproduced many observed features of pulsar glitches, they remain limited by the modest number of vortices and the small physical size of the simulated condensates; limitations that are unavoidable with current computational resources and that do not affect more phenomenological (non-microscopic) vortex-avalanche models. Building on this Gross–Pitaevskii framework, our study investigates the turbulent superfluid dynamics triggered by large vortex avalanches, uncovering a previously overlooked anomalous energy-injection mechanism.
@@
The inherent compressibility of BECs renders them susceptible to quantum turbulence (QT) under dynamic instabilities. This turbulence is primarily characterized by the breakdown of vortices in a self-similar process, leading to a power-law scaling of $k^{-5/3}$, known as the Kolmogorov spectrum [17, 18]. In addition to the Kolmogorov regime, QT in BECs also manifests in the Vinen regime, which is characterized by random vortex distributions in superfluids and exhibits a $k^{-1}$ scaling. Recent developments have identified novel turbulence regimes, including strong quantum turbulence [19, 20], marked by density fluctuations and non-polarized vortex lines, as well as rotational quantum turbulence [21, 22, 23, 24], and turbulence in self-gravitating dark-matter BEC candidates [25, 26].
@@
Previous studies have explored the nature of glitches and vortex avalanches in the context of pinning potentials and superfluid behavior. However, the role of superfluid-crust interactions in inducing turbulent flow remains largely unexamined. In this letter, we investigate the characteristics of turbulence and associated velocity flows arising from such interactions. Furthermore, it analyzes the mechanisms driving this turbulence, facilitated by pinning sites and the dynamic spin-down of the condensate. The damping effects on the system are also evaluated, identifying an optimal damping coefficient that maximizes turbulence strength. Recent studies indicate that quantum turbulence suppresses collective excitation modes in BECs [27, 28]. As these modes drive vortex avalanches and glitch phenomena, quantifying the onset of turbulence in such systems is essential. Beyond its implication for existing neutron star models, the effect of vortex depinning, particularly its role in driving incompressible turbulence through secondary injection, is of significant interest in the domain of quantum turbulence [29, 30, 31, 32, 19].
@@
where $\psi\equiv\psi(\mathbf{r},t)$ is the condensate wave function, with $\mathbf{r}\equiv(x,y)$. $\nabla^{2}=\partial_{x}^{2}\,+\partial_{y}^{2}$ represents the two-dimensional Laplacian, $\gamma$ corresponds to the parameterized damping coefficient, and $g$ is the nonlinear interaction strength given by $g=800$. Although the interaction parameter does not directly represent the microscopic nuclear interactions in a neutron superfluid, it models a short-range repulsive potential that captures the essential physics for generating quantum vortices. For the large-scale hydrodynamic behavior and collective vortex dynamics relevant to glitch-scale turbulence, the system can be described by effective theories where microscopic details are parameterized (e.g., into a mutual friction coefficient), rather than acting as the primary driver of the dynamics [34].
The time-dependent rotation frequency $\Omega(t)$ describes the rotational deceleration profile of the condensate as follows:
@@
where $\Omega_{0}$ is the initial rotation frequency of the condensate, and $t_{s}$ is the spin-down time. The $z$-component of the angular momentum is given by $L_{z}=\mathrm{i}\hbar(y\partial_{x}-x\partial_{y})$. In general, glitch events arise from transitions between two distinct non-zero rotational frequencies [35, 16]. However, a partial spin-down can retain vortices (see supplementary Fig.2), which may suppress the amplitude of glitches. While introducing a sudden deceleration can overcome this suppression, it often results in an unstable and unconfined condensate. Therefore, to enhance the vortex avalanche process without destabilizing the condensate, we have instead chosen to fully spin down the condensate.
@@
where $\tilde{V}(x_{\theta},y_{\theta})=3+\cos(n_{\mathrm{crust}}x_{\theta})+\cos(n_{\mathrm{crust}}y_{\theta})$, and the rotated coordinates are $x_{\theta}=\cos(\theta(t))x_{p}+\sin(\theta(t))y_{p}$ and $y_{\theta}=-\sin(\theta(t))x_{p}+\cos(\theta(t))y_{p}$. Here, $n_{\mathrm{crust}}$ determines the number of pinning sites, $r_{\mathrm{crust}}$ is the radius at which $V_{\mathrm{crust}}$ is maximized, and $\Delta r_{\mathrm{crust}}$ represents the crust thickness. The crust potential parameters are based on Refs. [15, 14]. The centrifugal potential, $V_{\mathrm{cent}}(\mathbf{r},t)=\frac{1}{2}\Omega(t)^{2}r^{2}$, ensures uniform condensate density under rotation. Following Refs. [33, 36, 37], we adopt a coherent length scale (comparable to vortex core size) of $\xi=10~\mathrm{fm}$, yielding an energy scale of $\epsilon=207~\mathrm{keV}$. With these parameters, the time scale is calculated as $\tau=3.2\times 10^{-21}~\mathrm{s}$. We compute the kinetic energy components and their spectra based on the analysis of their numerical implementation in Ref. [38]. It is important to note that while the GPE model is inherently limited to describing weakly interacting BECs in relatively small systems, resulting in a mismatch with the microscopic length scales present in neutron stars, it effectively captures vortex depinning phenomena, which are considered strong candidates for characterizing glitch events. Although scaling the vortex population, condensate radius and rotation frequency to astrophysical scales is unfeasible, the GPE model has reproduced Self-Organized Criticality of the pulsar glitch behavior  [16, 15] The simulation involves numerically solving the GPE using the split-step Crank-Nicolson method [39, 40] in a computational domain of size $512\times 512$, with spatial steps $dx=dy=0.05$ and a time step $dt=10^{-3}$ to ensure numerical stability and convergence. The numerical simulations are performed on an NVIDIA A100 GPU using CUDA C codes developed based on [41]. Initially, we prepare the condensate and crust with an initial rotation frequency $\Omega_{0}=2$ through imaginary-time iterations ($t\to-\mathrm{i}t$). This converged profile is then spun down at varying rates, characterized by the spin-down time $t_{s}$, in real-time. We first consider a case with a static crust and no dissipation ($\gamma=0$) while exploring the effects of damping in subsequent analyses.
@@
Further, the vortex avalanche triggered by Magnus flow results in an instantaneous recoupling of the superfluid and normal components, accompanied by a rapid transfer of angular momentum. This abrupt transition in superfluid systems may induce turbulence but cannot be confirmed using density profiles alone. To characterize the turbulent flow in the system, we therefore analyze the incompressible kinetic energy spectra of the condensate.
@@
As shown in Fig. 3, the vortex density exhibits a distinct scaling behavior with respect to time, further confirming the turbulent behavior of the condensate. During the initial Kolmogorov turbulent regime, the vortex-line density $l_{v}$ scales as $t^{-3/2}$, transitioning to $t^{-1}$ scaling in the later Vinen turbulent regime. This temporal decay of vortex-line density is attributed to vortex breakdown in the Kolmogorov regime, as well as to sound radiation (Kelvin-wave emission) and vortex reconnection in the Vinen regime [42, 43, 44]. For shorter spin-down times, the $t^{-3/2}$ scaling persists over a longer timescale, corresponding to a broader spatial extent of the $k^{-3/2}$ scaling in the incompressible kinetic energy spectrum (Fig. 2). Due to the intrinsic scales of the Gross–Pitaevskii framework, the power-law behavior observed here is limited to laboratory-scale systems. However, the functional form of this scaling, the $k^{-5/3}$ spectrum, is consistent with the incompressible turbulence spectrum reported in a distinct astrophysical context [25, 26]. This similarity in spectral behavior, despite the vast difference in physical scales and systems, suggests a common underlying turbulent mechanism may be at play.
@@
Although glitch events are explained by the transfer of angular momentum [45], analyzing the onset of turbulence and the dynamics of quantized vortices is required to fully understand the transport of kinetic energy components, which, in turn, enables a new perspective on vortex depinning. The evolution of net energy transport between the kinetic components, shown in Fig. 4, provides critical insight into the post-spin-down dynamics and explains the persistence of vortex flow. The energy exchange between the compressible–quantum pressure and incompressible–compressible components does not significantly influence the turbulent dynamics. However, the energy exchange between the incompressible and quantum pressure components, denoted $E_{\mathrm{ex}}^{iq}$, is initially negative, indicating energy transfer from the incompressible to the quantum pressure component. After the condensate spins down, the exchange becomes positive, indicating a transfer from quantum pressure to incompressible component and at later times $E_{\mathrm{ex}}^{iq}\to 0$, as no further incompressible energy is transferred to quantum pressure. For sufficiently short spin-down times [see Fig. 4(a)–(b)], the exchange energy becomes positive, indicating an injection of incompressible energy from the quantum pressure component. This secondary injection mechanism sustains turbulent behavior in the condensate, even after the loss of rotational forcing, which explains the existence of vortex flow after the spin-down time is reached. For longer spin-down times, the peak of the secondary injection diminishes, reducing the intensity of turbulence.
@@
The onset of turbulence in BECs with pinning potentials during spin-down provides crucial insight into the prevalent GPE models of neutron star cores [15, 14, 33]. The dependence of turbulence dynamics on spin-down profiles and dissipative interactions within our Gross-Pitaevskii analog allows us to map the model’s parameter space. This mapping is crucial for assessing the theoretical feasibility and inherent limitations of the analogy itself, thereby clarifying which aspects of neutron star glitch phenomenology it may, or may not, be suited to explore [33]. Furthermore, this study is valuable for exploring inhomogeneous two-dimensional quantum turbulence and its associated dynamics, particularly in relation to the emergence of secondary energy injection mechanisms. It is worth noting that in systems with a large number of vortices, secondary energy injection is primarily governed by the transfer of energy from compressible to incompressible modes. This process arises due to vortices traveling longer distances, repeatedly pinning and depinning at various locations. Similar to the partial spin-down scenario, this results in a substantial retention of vortices within the condensate, which inhibits the transfer of energy from quantum pressure to incompressible kinetic modes.
In conclusion, although the Gross-Pitaevskii analog necessarily simplifies the fermionic nature of neutron Cooper-pair superfluidity in neutron-star interiors—including density-dependent 1S0 and 3P2 pairing, possible hyperon contributions, and uncertainties in nuclear many-body physics—it remains the most microscopically faithful effective theory for studying the dynamics of individual quantized vortices and their mutual interactions. While directly applicable only to finite, laboratory-scale systems, it provides essential insight into the fundamental quantum mechanisms that may underpin vortex-mediated phenomena in astrophysical superfluids [33]. These additional physical contributions could, in principle, be incorporated through multi-component or binary-fluid extensions of the GP framework, which offers a promising direction for future research.
Our simulations reveal a previously unreported transient anomalous energy-injection mechanism, driven by quantum pressure, that operates during the highly turbulent phase immediately following major vortex avalanches. Although quantitative extrapolation to kilometre-scale neutron stars is not yet possible, and the effect diminishes under prolonged differential rotation, the consistent appearance of avalanche-like events, power-law glitch-size distributions, and sub-second rise times across numerous independent Gross-Pitaevskii studies supports the astrophysical relevance of the underlying vortex dynamics.
The anomalous energy-injection mechanism identified in the Gross–Pitaevskii analog system, therefore, represents a new and potentially relevant ingredient for future multi-scale and hybrid models of neutron star interiors.
@@ Acknowledgements.
A.S. acknowledges financial support from the Council of Scientific and Industrial Research (CSIR), India, in the form of a Direct Senior Research Fellowship. The work of P.M. is supported by the Ministry of Education-Rashtriya Uchchatar Shiksha Abhiyan (MoE RUSA 2.0): Bharathidasan University – Physical Sciences.